How many numbers between 000 and 999 (like a bike lock) contain the digit 9 at least once?
Answer #1

Total numbers = 1000

Total numbers which don't have digit 9 = 9 * 9 * 9 = 729

Total numbers with at least one digit 9 = 1000 - 729 = 271 (ans)